OpenStudy (anonymous):

interest = $8275 times 13% times 3 years

jimthompson5910 (jim_thompson5910):

interest = $8275 times 13% times 3 years interest = 8275*0.13*3 interest = 1075.75*3 interest = $3227.25

OpenStudy (anonymous):

yes thats what i got!